# Runtime Stealthy Perception Attacks against DNN-based Adaptive Cruise Control Systems

### v3

@@ V-D Comparison to DNN Output and Control Output Attacks
TABLE VIII: Performance of StrategicOut attack with all the safety features and different AEBS settings (AEBS with Shared Camera).
@@
All & AEBS Activated 100% 20.3% 80.1% (797/995)
All & AEBS Disabled 0.5% 99.5% 0.5%(5/1000)
We also evaluate the success rate of StrategicOut attack under two different safety intervention configurations. We do not assess the MaxOut attack due to its high likelihood of being detected. Table VIII shows that without AEBS, StrategicOut achieves a high success rate of 99.5% by generating attack values within safety limits and avoiding driver intervention. However, with AEBS active, using the same camera inputs as ACC, the success rate drops significantly to 20.3%, primarily due to safety interventions triggered in all simulations.
Figure 12: CA-Opt perception attack vs. DNN output attack.
Similarly, we compare the performance of CA-Opt attack with a stealthy attack that directly compromises DNN output 2 (referred to as DNNOut) by formulating an optimization problem to maximize the RD prediction within one standard deviation while ensuring the satisfaction of safety constraints on acceleration and speed [97]. We calculate the acceleration and speed values corresponding to RD predictions by replicating the Openpilot MPC and PID algorithms. This baseline uses the same context-aware method as CA-Opt for selecting the attack times and durations. We observe that the DNNOut attack causes a more obvious change in the RD predictions (see Fig. 12) compared to CA-Opt attack on DNN inputs 1, which then results in similar obvious changes in the gas, speed, or acceleration as depicted in Fig. 11.
@@
VI Evaluation in Real-World Settings
@@
RQ4: Is our attack robust to real-world factors such as different camera positions and weather conditions?
RQ5: Can our attack transfer well from simulation to real-world implementation?
RQ6: How does our attack affect the real-time system operation?
RQ7: Can our attack evade detection or mitigation by the existing adversarial patch defense methods?
VI-A Robustness to Real-world Factors
To assess attack robustness, we vary front camera height based on standard passenger car profiles from manufacturers [85]. We perform our experiments with four heights between 1.1-1.7 meters and three initial distances (50m, 75m, 100m).
Fig. 13 illustrates the 100% success rate of our CA-Opt attack across 12 testing scenarios. The Ego vehicle initially maintains a safe following distance, deviates from it around the 2,500-3,000 control cycle or step due to the adversarial patch, and eventually collides with the lead vehicle. These results demonstrate our attack is robust to different camera positions and initial longitudinal distances and can cause safety hazards.
We also test our attacks with diverse weather (rainy, sunny, or cloudy) and lighting conditions (noon or sunset). Results show that our CA-Opt attack causes longitudinal deviations of 9.8-14.3m in the predicted lead vehicle position, while maintaining a success rate of 100% under such conditions.
Figure 13: Actual relative distance trajectories under CA-Opt attack with different camera heights (H1:1.1m, H2:1.3m, H3:1.5m, H4:1.7m) and initial longitudinal distances to the lead vehicle (L1:50m, L2:75m, L3:100m). An actual relative distance of zero indicates collision.
VI-B Performance on Actual Vehicles
@@
Figure 14: Relative distance predictions with (solid lines) or w/o (dashed lines) adversarial patch for different driving scenarios.
@@
Then we conducted experiments using the same scenarios (SC1-SC4) outlined in Section V-A. For an accurate assessment of the impact of the attack, we cloned the LVD’s DNN model within the OpenPilot control software and ran both the original and the duplicate model on the AV simultaneously. During each perception cycle (20Hz), we initially supplied a benign image (with an odd image index number) to the standard DNN model. Then, we duplicated this benign image, injected the adversarial patch to it, and then fed it to the second DNN model. The predictions from each model were recorded in separate log files. The results, presented in Figure 14, show that the attack significantly increased the relative distance predictions for 15.3m on average in all tested scenarios.
@@
We also conducted experiments using a real-world video dataset as described in Appendix F. These experiments demonstrate the effectiveness of CA-Opt attack in impacting the DNN-based perception module in real-world driving scenarios.
@@
Figure 15: RD predictions with (1a) or w/o (1b) patch on an actual vehicle in a parking lot; (2a) Side view of lead car model; (2b) AV under perception attack collides with the lead car model; (2c) AV follows the car model in a benign scenario; (2d) Driver’s view upon collision. [Anonymous video: https://sites.google.com/view/CAP-Attack]
@@
End-to-End Evaluation. We also evaluate the impact of attacks on the end-to-end ACC on actual vehicle. To ensure the safety of both the driver and the vehicle, we constructed a lead car model from PVC pipe, designed to match the dimensions of a real BMW car model [16]. We aimed for the OpenPilot system to recognize this fabricated car as a genuine vehicle by attaching a rear-view image of a car to its rear end (see Fig. 15-2c). In this experiment, the AV approached the LV from a distance of 50 meters with a cruise speed set at 28 mph. Meanwhile, the LV was propelled by two remote-controlled ground robots (Fig. 15-2a). We conducted this experiment with and without the attack (adversarial patches) activated.
@@
OpenPilot software successfully recognized the lead car model as a legitimate vehicle and maintained a safe following distance and speed (about 5 mph) in the benign scenario (see Fig. 15-2c). However, in the presence of the attack, we observed that the AV continued to advance toward the lead car model and eventually collided with it (Fig. 15-2b), despite the AEBS being activated (Fig. 15-2d). This underscores the generalization of our proposed attack in efficiently causing safety hazards and exposes the inadequacy of existing safety mechanisms in preventing the attack. Moreover, the time elapsed between the AEBS warning and the collision was approximately 1.5-2.2 seconds, shorter than the average driver reaction time of 2.5 seconds, leaving insufficient time for a human driver to intervene and prevent the collision.
VI-C Runtime Overhead
Experimental results show that the time overhead introduced by the context inference component before activating the attacks is minimal (1.17 us). Following the activation of attacks, the time overhead for the object detection module is about 10.1 ms on average. Note that some production ADAS provide object detection and tracking features, so this overhead time could be potentially avoided. We also observe that the primary attribution algorithm [83] does not add significant overhead, leveraging gradients calculated during the patch optimization process. The total time overhead is 1.52 ms.
Figure 16: Runtime overhead of each step of the attack.
VI-D Evading Existing Defense Methods
@@
While the proposed CA-Opt attack can create stealthy adversarial patches invisible to the human eye, it may be detected by some existing defense methods.
@@
Adversarial Patch Detection. Methods such as gradient masking [69], lossy compression [49], or adversarial training [61] have been proposed for adversarial patch detection. However, these methods either need to be trained on specific attacks with high computation costs [92, 27, 54] or significantly sacrifice the DNN prediction accuracy [69, 91], which could negatively affect the safety of ACC systems.
@@
We assess four widely used open-source defense methods that only rely on model input transformation without the need for re-training, including adding Gaussian noise [94], JPEG compression [33], reducing image color bit-depth [93], and using spatial median smoothing [93]. We evaluate the attack success rates in causing hazards under each defense method with various parameter settings while considering the effect of input transformations on the benign or attack-free image frames to maintain the baseline safety of the ACC system.
Figure 17: Results of each directly-applicable defense method.
@@
As shown in Fig. 17, JPEG compression and bit-depth reduction methods effectively reduce the attack success rate in causing hazards under specific parameter configurations. However, these methods fall short in maintaining the ACC’s safety by leading the benign image frames to cause hazards. In instances where the benign cases do not lead to hazards, the attack hazard rate is at 100%. On the other hand, the incorporation of Gaussian noise or median smoothing reduces the ACC’s LV detection accuracy. These methods are ineffective in mitigating CA-Opt attacks (hazard rate stays at 100% for all configurations), while also causing hazards for benign frames.
@@
Sensor Fusion. An alternative defense against adversarial patches could involve integrating independent sensors like Lidar or radar with camera data for LVD predictions. However, Lidar is too costly for Level-2 AVs [76], and our tests found that radar-camera fusion did not prevent ACC misbehavior or collisions (see Appendix A). This may be because of the use of Kalman filters in sensor fusion, which assumes measurement noise is zero-mean Gaussian and are vulnerable to perturbations smaller than one standard deviation of this noise [44]. In addition, sensor fusion outputs a weighted summation of radar and camera predictions, which cannot completely eliminate deviations caused by erroneous camera predictions, particularly when they significantly deviate from the ground truth. In some production ACC, camera predictions typically carry more weight. The vulnerability of sensor fusion was also reported in previous works [40]
VII Discussion
@@
Sim-to-Real Gap. Addressing the sim-to-real gap in AV security literature is challenging due to the risks and costs of real-road tests. In this paper, we tried to narrow this gap by developing a realistic experimental platform that integrates production ADAS control software, a physical-world simulator, and well-designed safety interventions with high-risk driving scenarios designed based on the NHTSA report [67]. Moreover, we evaluate the sim-to-real transfer possibility using an actual vehicle, a model lead car, and a publicly available dataset. However, limitations exist, such as the fixed model and thresholds used for the human driver simulator that may impact evaluation results. Also, additional firmware and safety checks might be deployed in real cars, which can further limit the effectiveness of the proposed attack.
@@
Attack Method Generalization. We demonstrate the generalization of our proposed attack on a production ACC system, OpenPilot, through closed-loop simulation, real-world AV dataset, and actual vehicle experiments. However, the vulnerability of other Level-2 production ACC systems, such as Tesla Autopilot or Cadillac Super Cruise, to our attacks remains uncertain due to their closed-source nature. While we cannot directly evaluate our attacks on these systems, it is reasonable to argue that our results hold generalization potential based on the representative nature of the OpenPilot ACC system. Specifically, our attack strategy, which leverages context awareness derived from high-level system hazard analysis, is not limited to specific systems and can be generalized to diverse ACC systems. Furthermore, our optimization-based attack vector generation can be applied to other DNN-based ACC systems, given the inherent vulnerability of DNNs to adversarial input perturbations [36, 76, 44, 37].
VIII Related Work
@@
Adversarial Attacks on DNN. Many works have explored the vulnerability of DNN against adversarial attacks by adding adversarial physical/digital patches or stickers [55, 51, 41, 37, 65, 36, 76, 84, 90]. However, most of these works focus on altering the prediction class or probability or lane line position, which do not apply to attacks against ACC. Moreover, they rely on off-line optimization of attack value, neglecting the impact of attack timing. In contrast, our work introduces a novel runtime perception attack method against DNN-based ACC systems, employing a combined knowledge and data-driven approach that considers both attack timing and value for enhanced effectiveness. The only other work on ACC [39] focused on the physical attacks without considering dynamic changes at runtime, which is not scalable to many vehicles.
@@
Security Analysis of AVs. Great efforts have also been made in studying the security of AVs, such as the security of Lidar [75], GPS [81], radar [47], camera [74], lane detection [76, 84], multiple objects tracking [44, 45, 58], control software [97], and safety mechanisms [60]. To the best of our knowledge, this paper is the first analysis of the security of Level-2 production ACC systems under stealthy safety-critical attack by considering three levels of safety interventions by constraint checking, human driver, and AEB/FCW and addressing unique challenges (Section III-B).
IX Conclusion
@@ Appendix A Sensor Fusion
We implement a radar sensor in the CARLA simulator and feed the data to the OpenPilot radar interface [88], to be used as an independent input by the fusion module. Specifically, we use the DBSCAN algorithm [79] to cluster the 2D point map of the relative distance and speed of perceived objects from the radar sensor and feed their mean values to OpenPilot, which are then further filtered and processed for fusion.
Fig. 18 (Top) shows an example of predictions of the relative distance to the lead vehicle from the fusion of the camera and radar measurements. We see that the radar and camera predictions agree well most of the time. Also, the error between the fusion predictions and the ground truth relative distance (based on positions of vehicles in the simulator) becomes smaller as the Ego vehicle approaches the lead vehicle (an RMSE of 0.81m after 3000 control cycles in the figure). Sensor fusion also helps reduce the errors in fusion predictions under attacks as shown in Fig. 18 (Bottom) and discussed in Section VI-D, even though it fails to prevent collisions in the end.
Figure 18: An example fusion of relative distance predictions based on camera and radar data compared with the ground truth under normal operation (Top) or under attack (Bottom).

### v4

@@ 5.4. Comparison to DNN Output and Control Output Attacks
We also evaluate the success rate of StrategicOut attack under two different safety intervention configurations. We do not assess the MaxOut attack due to its high likelihood of being detected. Table 8 shows that without AEBS, StrategicOut achieves a success rate of 81.9% by generating attack values within safety limits and avoiding driver intervention. However, with AEBS active, using the same camera inputs as ACC, the success rate drops significantly to 20.3%, due to AEBS interventions triggered in all simulations.
Table 8. Performance of StrategicOut attack with all the safety features and different AEBS settings (AEBS with Shared Camera).
@@
All & AEBS Activated StrategicOut 20.3% 79.7% (797/1,000)
All & AEBS Disabled StrategicOut 81.9% 18.1%(181/1,000)
All & AEBS Activated OptOut 34.5% 65.5 (655/1,000)
Figure 11. Context-Aware perception attacks vs. output attacks.
We further compared the CA-Opt attack with a stealthy control output attack that causes the exact deviations of the state variables as the proposed perception attack (referred to as OptOut). Specifically, we reran the simulations and injected the faults by setting the control output to the recorded output traces caused by the CA-Opt perception attack. We observed that the OptOut attack achieved a higher success rate (34.5%) than the StrategicOut attack. However, it did not change the DNN predictions or affect the AEBS function, thus triggering safety interventions more easily and earlier than the CA-Opt attack. In addition, CAN outputs are encrypted in some car models (jha2020ml, ), increasing the attack cost.
DNN Output Attacks. Similarly, we compare CA-Opt attack with a stealthy attack that directly compromises DNN output 2 (referred to as DNNOut) by formulating an optimization problem to maximize the RD prediction within one standard deviation while ensuring the satisfaction of safety constraints on acceleration and speed (zhou2022strategic, ). We calculate the acceleration and speed values corresponding to RD predictions by replicating the Openpilot MPC and PID algorithms. This baseline uses the same context-aware method as CA-Opt for selecting the attack times and durations. We observe that the DNNOut attack causes a more obvious change in the RD predictions (see Fig. 12) compared to CA-Opt attack on DNN inputs 1, which then results in similar obvious changes in the gas, speed, or acceleration as depicted in Fig. 11.
@@
Figure 12. CA-Opt perception attack vs. DNN output attack.
6. Evaluation in Real-World Settings
@@
RQ4: Can our attack transfer well from simulation to real-world implementation?
RQ5: Can our attack evade detection or mitigation by the existing adversarial patch defense methods?
We also have evaluated the runtime overhead of our attack and its robustness to real-world factors, as described in Appendix F-G.
6.1. Performance on Actual Vehicles
@@
Figure 13. Relative distance predictions with (solid lines) or w/o (dashed lines) adversarial patch for different driving scenarios.
@@
Then, we conducted experiments using the same scenarios (SC1-SC4) outlined in Section 5.1. For an accurate assessment of the impact of the attack, we cloned the LVD’s DNN model within the OpenPilot control software and ran both the original and the duplicate model on the AV simultaneously. During each perception cycle (20Hz), we initially supplied a benign image to the standard DNN model. Then, we duplicated this benign image, injected the adversarial patch to it, and then fed it to the second DNN model. The predictions from each model were recorded in separate log files. The results, presented in Fig. 13, indicate that the attack increased RD predictions by an average of 15.3 meters across all tested scenarios.
@@
We also conducted experiments using a real-world video dataset as described in Appendix H. These experiments demonstrate the effectiveness of CA-Opt attack in impacting the DNN-based perception module in real-world driving scenarios.
@@
Figure 14. RD predictions w/o (1a) or w (1b) patch on an actual vehicle in a parking lot; (2a) Side view of lead car model; (2b) AV under perception attack collides with the lead car model; (2c) AV follows the car model in a benign scenario; (2d) Driver’s view upon collision.
@@
End-to-End Evaluation. We also evaluate the impact of attacks on the end-to-end ACC on actual vehicle. To ensure the safety of both the driver and the vehicle, we constructed a lead car model from PVC pipe, designed to match the dimensions of a real BMW car model (bmwsize, ). We aimed for the OpenPilot system to recognize this fabricated car as a genuine vehicle by attaching a rear-view image of a car to its rear end (see Fig. 14-2c). In this experiment, the AV approached the LV from a distance of 50 meters with a cruise speed set at 28 mph. Meanwhile, the LV was propelled by two remote-controlled ground robots (Fig. 14-2a). We conducted this experiment with and without attack (adversarial patches) activated.
@@
OpenPilot software successfully recognized the lead car model as a legitimate vehicle and maintained a safe following distance and speed (about 5 mph) in the benign scenario (see Fig. 14-2c). However, in the presence of the attack, the AV continued to advance toward the lead car and eventually collided with it (Fig. 14-2b), despite AEBS being activated (Fig. 14-2d). This underscores the generalization of our proposed attack in efficiently causing safety hazards and exposes the inadequacy of existing safety mechanisms in preventing the attack. Moreover, the time between the AEBS warning and the collision was approximately 1.5-2.2 seconds, shorter than the average driver reaction time of 2.5 seconds, leaving insufficient time for a human driver to intervene and prevent the collision.
6.2. Evading Existing Defense Methods
@@
While our proposed stealthy adversarial patches are invisible to human eye, they may be detected by some existing defense methods.
@@
Adversarial Patch Detection. Methods such as gradient masking (papernot2016towards, ), lossy compression (kurakin2018adversarial, ), or adversarial training (madry2017towards, ) have been proposed for adversarial patch detection. However, these methods either need to be trained on specific attacks with high computation costs (xu_patchzero_2023, ; chen_jujutsu_2023, ; liu_segment_2022, ) or significantly sacrifice DNN prediction accuracy (papernot2016towards, ; xiang_patchguard_nodate, ), which negatively affect ACC systems’ safety.
@@
We assess four widely used open-source defense methods that only rely on model input transformation without the need for re-training, including adding Gaussian noise (zhang2019defending, ), JPEG compression (dziugaite2016study, ), reducing image color bit-depth (xu2017feature, ), and using spatial median smoothing (xu2017feature, ). We implement all these defense methods by changing each input image frame with various parameter settings (as shown in the x-axis of Fig. 15). We evaluate the attack success rates in causing hazards under each defense method while considering the effect of input transformations on the benign or attack-free image frames to maintain the baseline safety of the ACC system.
Figure 15. Results of each directly-applicable defense method.
@@
As shown in Fig. 15, JPEG compression and bit-depth reduction methods effectively reduce the attack success rate under specific parameter configurations. However, these methods fall short in maintaining the ACC’s safety by leading the benign image frames to cause hazards. In instances where the benign cases do not lead to hazards, the attack hazard rate is at 100%. On the other hand, the incorporation of Gaussian noise or median smoothing reduces the ACC’s LV detection accuracy. These methods are ineffective in mitigating CA-Opt attacks (hazard rate stays at 100% for all configurations), while also causing hazards for benign frames.
These results indicate that our attack can easily evade the directly applicable defense methods. More advanced methods, such as adversarial training, may need to be developed/trained specifically for our design, which are subject to future direction.
@@
Sensor Fusion. An alternative defense against adversarial patches could involve integrating independent sensors like Lidar or radar with camera data for LVD predictions. However, Lidar is too costly for Level-2 AVs (sato2021dirty, ), and our tests found that radar-camera fusion did not prevent ACC misbehavior or collisions (see Appendix A). This may be because of using Kalman filters for sensor fusion, which assume measurement noise is zero-mean Gaussian and are vulnerable to perturbations smaller than one standard deviation of this noise (jha2020ml, ). In addition, sensor fusion outputs a weighted summation of radar and camera predictions, which cannot completely eliminate deviations caused by erroneous camera predictions, particularly when they significantly deviate from the ground truth. In some production ACC, camera predictions carry more weight. The sensor fusion vulnerability was also reported in previous work (hallyburton2022security, ).
7. Discussion
@@
Sim-to-Real Gap. Addressing the sim-to-real gap in AV security literature is challenging due to the risks and costs of real-road tests. In this paper, we tried to narrow this gap by developing a realistic experimental platform that integrates production ADAS control software, a physical-world simulator, and well-designed safety interventions with high-risk driving scenarios designed based on the NHTSA report (najm_pre-crash_2007, ). Moreover, we evaluate the sim-to-real transfer possibility using an actual vehicle, a model lead car, and a publicly available dataset. However, there are still some limitations, such as using fixed models and thresholds in the design of the driver reaction simulator, that may impact evaluation results. Also, additional firmware and safety checks might be deployed in real cars, which can further limit the effectiveness of the proposed attack.
@@
Attack Method Generalization. We demonstrate the generalization of our proposed attack on a production ACC system, OpenPilot, through closed-loop simulation, real-world AV dataset, and actual vehicle experiments. However, the vulnerability of other Level-2 production ACC systems, such as Tesla Autopilot or Cadillac Super Cruise, to our attacks remains uncertain due to their closed-source nature. While we cannot directly evaluate our attacks on these systems, it is reasonable to argue that our results hold generalization potential based on the representative nature of the OpenPilot ACC system. Specifically, our attack strategy, which leverages context awareness derived from high-level system hazard analysis, can be generalized to diverse ACC systems. Furthermore, our optimization-based attack vector generation can be applied to other DNN-based ACC systems, given the inherent vulnerability of DNNs to adversarial input perturbations (eykholt2018robust, ; sato2021dirty, ; jha2020ml, ; goodfellow2014explaining, ).
8. Related Work
@@
Adversarial Attacks on DNN. Many works have explored the vulnerability of DNN against adversarial attacks by adding adversarial physical/digital patches or stickers (liu2018dpatch, ; lee2019physical, ; hoory2020dynamic, ; goodfellow2014explaining, ; moosavidezfooli2016deepfool, ; eykholt2018robust, ; sato2021dirty, ; tencent2019experimental, ; wu2020making, ). However, most of these works focus on altering the prediction class or probability or lane line position, which do not apply to attacks against ACC. Moreover, they rely on off-line optimization of attack value, neglecting the impact of attack timing. In contrast, our work introduces a novel runtime perception attack method against DNN-based ACC systems, employing a combined knowledge and data-driven approach that considers both attack timing and value for enhanced effectiveness. The only other work on ACC (guo2023adversarial, ) focused on the physical attacks without considering dynamic changes at runtime, which is not scalable to many vehicles.
@@
Security Analysis of AVs. Great efforts have also been made in studying the security of AVs, such as the security of Lidar (lidar_2023, ), GPS (shen2020drift, ), radar (komissarov2021spoofing, ), camera (sato_wip_2023, ), lane detection (sato2021dirty, ; tencent2019experimental, ), multiple objects tracking (jha2020ml, ; jia2019fooling, ; ma_wip_2023, ), control software (zhou2022strategic, ; ding2021mini, ), and safety mechanisms (ma2021sequential, ). To the best of our knowledge, this paper is the first analysis of the security of Level-2 production ACC systems under stealthy safety-critical attack by considering three levels of safety interventions by constraint checking, human driver, and AEB/FCW and addressing unique challenges (Section 3.2).
9. Conclusion
@@
Acknowledgment
This work was partially supported by a gift from Toyota InfoTechnology Center and by the National Science Foundation (NSF) under Grants 2402941 and 1931997.
@@ Appendix A Sensor Fusion
We implement a radar sensor in the CARLA simulator and feed the data to the OpenPilot radar interface (weng2022formal, ), to be used as an independent input by the fusion module. Specifically, we use the DBSCAN algorithm (schubert2017dbscan, ) to cluster the 2D point map of the relative distance and speed of perceived objects from the radar sensor and feed their mean values to OpenPilot, which are then further filtered and processed for fusion.
Fig. 16 (Top) shows an example of predictions of the relative distance to the lead vehicle from the fusion of the camera and radar measurements. We see that the radar and camera predictions agree well most of the time. Also, the error between the fusion predictions and the ground truth relative distance (based on positions of vehicles in the simulator) becomes smaller as the Ego vehicle approaches the lead vehicle (an RMSE of 0.81m after 3,000 control cycles in the figure). Sensor fusion also helps reduce the errors in fusion predictions under attacks as shown in Fig. 16 (Bottom) and discussed in Section 6.2, even though it fails to prevent collisions in the end.
Figure 16. An example fusion of relative distance predictions based on camera and radar data compared with the ground truth under normal operation (Top) or under attack (Bottom).